a. What would be the gravitational force exerted on a 65.0 kg person on Jupiter? Jupiter mass is 1.90 X 1027 kg and it s diameter is 1.43 X 108 meters?

b. What is the gravitational force of 65 kg person on Earth?

c. What is the ratio of a 65 kg person weight on Jupiter and Earth?
Answer #1

here,

a)

mass of person , m = 65 kg

the mass of Jupiter , M1 = 1.9 * 10^27 kg

the diameter of Jupiter , d = 1.43 * 10^8 m

the radius of Jupiter , r = d/2 = 7.15 * 10^7 m

the gravitational force on 65 kg person , W1 = G * M1 * m /r^2

W1 = 6.67 * 10^-11 * 1.9 * 10^27 * 65 /(1.43 * 10^8)^2 N

W1 = 403.8 N

b)

the weight on Earth , W2 = m * g

W2 = 65 * 9.81 N = 637.65 N

c)

the ratio of a 65 kg person weight on Jupiter and Earth , R = W1 /W2

R = 403.8 /637.65 = 0.63